Explain, in your own words, how to evaluate a polynomial for a given value of the variable. Demonstrate the process with an example.

Create a polynomial for the class to evaluate for two different given values.

Answer by jainpreeja(10)

put the value in the variable place and then solve it
Example:f(x)=2x^2-3x+6 find the value where x=3
f(3)=2*(3)^2-3*3+6
=2*9-9+6
=18-9+6
=9+6
=15